OpenStudy (anonymous):

ANSWER AND I'LL FAN YOU! The table shows the total number of cupcakes y Lucas plans to buy if x guests are coming to his birthday party. Using the function rule shown by the table, what will the dependent variable be when the independent variable is twelve? A table of x and y-values. When x is 1, y is 5. When x is 4, y is 8. When x is 7, y is eleven. When x is ten, y is fourteen. 8 9 sixteen seventeen

OpenStudy (anonymous):

The independent variable is the x value, the dependent variable is the y value. You can derive the function used from the values in the table. From the table we can see that the change in x is constant ie it is 3, the change in y is also constant at 3. This defines a linear function with gradient =1 since change in x =change in y. To find the y intercept , derive the equation of the line using y=mx+c where m=1 and (x,y) is any point from the given points. c=4 therefore when x=12, y=12+4=16.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

pleasure